# KDP as a thermal blocking filter Deep near IR observations with a warm narrow band filter

### v1

@@ 4.1 KDP transmission and optical density
4.2 KDP equilibrium temperature and operating point
@@
where is the total radiant power (Eq. 4), =136.086 molecular weight of KDP, =2.338 g cm the density of KDP (W. M. Haynes et al., 2016), and the molar heat capacity as tabulated by C. C. Stephenson & J. G. Hooley (1944). Assuming the emissivity to be unity, we find that it takes about 20–35 min to warm the 0.5 thermally insulated KDP above its , and make the unwanted structure to disappear (Fig. 8). Given =288 K we find the KDP to reach an equilibrium temperature of 150 K when left in the beam.
4.3 Data reduction
Observations consisted of either 5-point or 9-point skewed dither-patterns with 10–15″ offsets between two consecutive integrations. All data were recorded as ramp-sampling integrations (Sec. 2.4). The broadband standard star fields were recorded with a sample time of 4 s with a number of samples yielding a total integrated flux within the 1% linearity range of the detector. The NB integrations had very low flux levels and were readout with the maximum 14 ramp-samples available. All sub-reads are stored, and we re-fit to the sub-reads ignoring the first read in a ramp-sample sequence. In the NOTCam data, the first line and row of each detector quadrant reads out as value 0. These together with hot and cold pixels were removed by bilinear interpolation based on the surrounding pixel values.
The images were then sky subtracted, aligned, and co-added with ’notcam.cl’ reduction package (priv. comm. A. A. Djupvik, T. Reynolds)222www.not.iac.es/instruments/notcam written for Image Reduction and Analysis Facility (IRAF) (D. Tody, 1986, 1993). The first image recorded for each dither-pattern was rejected from the co-addition. For the sky level being very low in the narrowband images, the NB images were additionally co-added without sky subtraction for estimating the sky level, and the instrumental background in case of KDP observations. Isophotal fluxes of the standards stars were extracted, and extinction corrected using the standard extinction coefficients adopted in the NOT’s signal-to-noise calculator extinction co-efficients (, , , , mag arcsec-2). We derive the zeropoints in units of ADU s-1 from the photometric nights (January 14, January 15, and June 30), and apply them to the sky surface brightness measurement of all nights. We have not derived color terms for the BP1191-4 filter, and use the J-band magnitudes for relative photometry in the BP1191-4 bandpass.
4.4 Instrumental background
@@
4.5 Thermal emission from the KDP
@@
4.6 The effect of persistence and reciprocity failure
We do not have a sophisticated model available for correcting persistence and reciprocity failure (S. Tulloch et al., 2019). We do see a power law decay of the background level, in the non-illuminated detector areas in our KDP integrations. We attribute the exponential decay to persistence, and subtract it out as a parasitic signal. Additionally, in case of the long BP1191-4 integrations we see spatial correlation in our sky images with the location of the brightest stars. The full field-of-view in the observations without KDP allows us to find a sky area which has not been visited by a star in the last few frames. However, in the integrations with the KDP, the usable image area is significantly smaller (Fig. 6), restricting the sky are to be measured in the proximity of the field stars. If the effect in the BP1191-4 observations with and without KDP is similar, this could lead us to over estimate the sky surface brightness by 15% in the narrowband KDP observations.
4.7 Airglow continuum at 1.191
@@ 4.6 The effect of persistence and reciprocity failure
To isolate the atmospheric continuum emission background, we calculate the zodiacal light brightness for our pointings with the InfraRed Science Archive (IRSA), Infrared Processing & Analysis Center (IPAC) Euclid background model calculator Version 1333https://irsa.ipac.caltech.edu/applications/BackgroundModel/. The calculator is based on the zodiacal light model of T. Kelsall et al. (1998). Both of the standard star fields observed in the BP1191-4 bandpass are located approximately at ecliptic latitude of =45, and the zodiacal light contributes only 35–40 , or 21.8 mag arcsec-2, which is much less than the sky surface brightness we measure.
In order to estimate the off-band sky light not rejected by our filter setup, and the potential contribution from emission lines in the wings of the BP1191-4 bandpass, we take the sky radiance model of ESO SkyCalc (S. Noll et al., 2012; A. Jones et al., 2013) for La Silla. We do not know the exact quantum efficiency figure for our detector but assume it to be 65% in the range of 0.5–2.6 . Assuming 95% reflectivity on the telescope mirrors, 99% transmission on all lenses, and 7.2% Fresnel losses on the KDP window, we estimate the sky lines in the wings of the BP1191-4 filter bandpass to contribute 9ADU s-1 px-1 to the measured sky surface brightness. Likewise, the integrated sky emission outside of the BP1191-4 bandpass would contribute 1ADU s-1 px-1. The transmission and reflectivity figures are optimistic and should give an upper limit on the excess sky contribution. We conclude that we do not see significant amount of unwanted light in our measurement.
H. T. Nguyen et al. (2016) has measured 18.8 mag arcsec-2 in a bandpass similar to ours at the Table Mountain Observatory, Big Pines, California. Other narrowband measurements in a similar bandpass, have measured a brighter sky surface brightness of 17.5 mag arcsec-2 (see B. Milvang-Jensen et al., 2013, for a compilation). Spectroscopic observations of P. W. Sullivan & R. A. Simcoe (2012) at Las Campanas, Chile indicate similar values, with a median dark time sky continuum brightness of 18.6 mag arcsec-2. S. Noll et al. (2024) gives a yearly average continuum background of 230 , or 19.8 mag arcsec-2 at 1.191 for Cerro Paranal, Chile. The differences might be explainable by the climatological variability of the hydroperoxyl continuum emission (S. Noll et al., 2025).
@@
5 Discussion
5.1 KDP as a thermal blocking filter
@@
5.2 Possibilities in filter design
KDP and its isomorphisms open up new possibilities for NIR filter design since they can be used as a shortpass filter ’glass’. We have designed and demonstrated use of a highly efficient narrowband filter operating outside the instrument cryostat (Fig. 3). Furthermore, we show a design for a highly optimized ’blue’ NIR filter to be used with a 2.5 cutoff HgCdTe detector. The implementation of the Z band (UKIDSS Z444http://www.ukidss.org/technical/instrument/filters.html, P. C. Hewett et al. (2006)) allows high peak transmission, and well defined bandpass edges due to the relaxed blocking requirements of the dielectric coating at wavelengths 1.9 , and high off-band blocking density due to KDP (Fig. 4).
5.3 Alternative materials
@@
5.4 Possible caveats
KDP contains Potassium which has radioactive isotopes. The radioactive decay leads to particle events observable by the photo-detectors if not shielded by a cover glass. KDP is a birefringent material, and the cutoff wavelength has a polarization dependence, which may affect some use cases. KDP is also water soluble (M. J. Weber, 1994), and a certain amount of care should be taken handling it.
6 Conclusion
We have shown that the KDP is an excellent thermal blocking filter and practical to use in an astronomical context, with limitation on its operating temperature, and birefringence. KDP offers good optical characteristics, and can strongly reduce atmospheric or instrumental thermal background. We have demonstrated the use of a warm narrow band filter in conjunction with a 2.5 cutoff detector, which has been enabled by the high thermal IR absorbance of the KDP. While we cannot demonstrate a photometric gain in the warm narrowband over the J band, we have measured the airglow continuum brightness within the bandpass, and find it to be 18.5 mag arcsec-2. KDP has the potential to become a mainstay in the SWIR astronomy as a heat absorbing, thermal emission blocking, shortpass filter material. It offers a way to improve the SWIR broad- and especially narrowband filter efficiency, and off-band blocking characteristics, can be used for enhancing performance of existing instrumentation suffering from thermal leaks, and has potential for realizing new instrument concepts.

### v2

@@ IV.1 KDP transmission and optical density
IV.2 KDP equilibrium temperature and operating point
@@
where is the total radiant power (Eq. 4), =136.086 molecular weight of KDP, =2.338 g cm the density of KDP (CRCchem), and the molar heat capacity as tabulated by Stephenson44a. Assuming the emissivity to be unity, we find that it takes about 20–35 min to warm the 0.5 thermally insulated KDP above its , and make the unwanted structure to disappear (Fig. 8). Given =288 K we find the KDP to reach an equilibrium temperature of 150 K when left in the beam.
IV.3 Data reduction
Observations consisted of either 5-point or 9-point skewed dither-patterns with 10–15″ offsets between two consecutive integrations. All data were recorded as ramp-sampling integrations (Sec. II.4). The broadband standard star fields were recorded with a sample time of 4 s with a number of samples yielding a total integrated flux within the 1% linearity range of the detector. The NB integrations had very low flux levels and were readout with the maximum 14 ramp-samples available. All sub-reads are stored, and we re-fit to the sub-reads ignoring the first read in a ramp-sample sequence. In the NOTCam data, the first line and row of each detector quadrant reads out as value 0. These together with hot and cold pixels were removed by bilinear interpolation based on the surrounding pixel values.
The images were then sky subtracted, aligned, and co-added with ’notcam.cl’ reduction package (priv. comm. A. A. Djupvik, T. Reynolds)222www.not.iac.es/instruments/notcam written for Image Reduction and Analysis Facility (IRAF) (IRAF1; IRAF2). The first image recorded for each dither-pattern was rejected from the co-addition. For the sky level being very low in the narrowband images, the NB images were additionally co-added without sky subtraction for estimating the sky level, and the instrumental background in case of KDP observations. Isophotal fluxes of the standards stars were extracted, and extinction corrected using the standard extinction coefficients adopted in the NOT’s signal-to-noise calculator extinction co-efficients (, , , , mag arcsec-2). We derive the zeropoints in units of ADU s-1 from the photometric nights (January 14, January 15, and June 30), and apply them to the sky surface brightness measurement of all nights. We have not derived color terms for the BP1191-4 filter, and use the J-band magnitudes for relative photometry in the BP1191-4 bandpass.
IV.4 Instrumental background
@@
IV.5 Thermal emission from the KDP
@@
IV.6 The effect of persistence and reciprocity failure
We do not have a sophisticated model available for correcting persistence and reciprocity failure (Tulloch19). We do see a power law decay of the background level, in the non-illuminated detector areas in our KDP integrations. We attribute the exponential decay to persistence, and subtract it out as a parasitic signal. Additionally, in case of the long BP1191-4 integrations we see spatial correlation in our sky images with the location of the brightest stars. The full field-of-view in the observations without KDP allows us to find a sky area which has not been visited by a star in the last few frames. However, in the integrations with the KDP, the usable image area is significantly smaller (Fig. 6), restricting the sky are to be measured in the proximity of the field stars. If the effect in the BP1191-4 observations with and without KDP is similar, this could lead us to over estimate the sky surface brightness by 15% in the narrowband KDP observations.
IV.7 Airglow surface brightness at 1.191
@@ IV.6 The effect of persistence and reciprocity failure
To isolate the atmospheric continuum emission background, we calculate the zodiacal light brightness for our pointings with the InfraRed Science Archive (IRSA), Infrared Processing & Analysis Center (IPAC) Euclid background model calculator Version 1333https://irsa.ipac.caltech.edu/applications/BackgroundModel/. The calculator is based on the zodiacal light model of Kelsall98. Both of the standard star fields observed in the BP1191-4 bandpass are located approximately at ecliptic latitude of =45, and the zodiacal light contributes only 35–40 , or 21.8 mag arcsec-2, which is much less than the sky surface brightness we measure.
In order to estimate the off-band sky light not rejected by our filter setup, and the potential contribution from emission lines in the wings of the BP1191-4 bandpass, we take the sky radiance model of ESO SkyCalc (Noll12; Jones13) for La Silla. We do not know the exact quantum efficiency figure for our detector but assume it to be 65% in the range of 0.5–2.6 . Assuming 95% reflectivity on the telescope mirrors, 99% transmission on all lenses, and 7.2% Fresnel losses on the KDP window, we estimate the sky lines in the wings of the BP1191-4 filter bandpass to contribute 9ADU s-1 px-1 to the measured sky surface brightness. Likewise, the integrated sky emission outside of the BP1191-4 bandpass would contribute 1ADU s-1 px-1. The transmission and reflectivity figures are optimistic and should give an upper limit on the excess sky contribution. We conclude that we do not see significant amount of unwanted light in our measurement.
Nguyen16 has measured 18.8 mag arcsec-2 in a bandpass similar to ours at the Table Mountain Observatory, Big Pines, California. Other narrowband measurements in a similar bandpass, have measured a brighter sky surface brightness of 17.5 mag arcsec-2 (see Milvang13, for a compilation). Spectroscopic observations of Sullivan12 at Las Campanas, Chile indicate similar values, with a median dark time sky continuum brightness of 18.6 mag arcsec-2. Noll24 gives a yearly average continuum background of 230 , or 19.8 mag arcsec-2 at 1.191 for Cerro Paranal, Chile.
@@
V Discussion
V.1 KDP as a thermal blocking filter
@@
V.2 Possibilities in filter design
KDP and its isomorphisms open up new possibilities for NIR filter design since they can be used as a shortpass filter ’glass’. We have designed and demonstrated use of a highly efficient narrowband filter operating outside the instrument cryostat (Fig. 3). Furthermore, we show a design for a highly optimized ’blue’ NIR filter to be used with a 2.5 cutoff HgCdTe detector. The implementation of the Z band (UKIDSS Z444http://www.ukidss.org/technical/instrument/filters.html, Hewett06) allows high peak transmission, and well defined bandpass edges due to the relaxed blocking requirements of the dielectric coating at wavelengths 1.9 , and high off-band blocking density due to KDP (Fig. 4).
V.3 Alternative materials
@@
V.4 Possible caveats
KDP contains Potassium which has radioactive isotopes. The radioactive decay leads to particle events observable by the photo-detectors if not shielded by a cover glass. KDP is a birefringent material, and the cutoff wavelength has a polarization dependence, which may affect some use cases. KDP is also water soluble (CRCoptical), and care should be taken handling it. Since we had no prior experience with the material, we kept the KDPs inside their vacuum bags until the measurement and installation in the instrument. We used standard nitrile gloves and face masks when handling the crystals with the expectation that the moisture from hands or breath could be sufficient to damage the uncoated optical surfaces. The instrument was closed and evacuated from air soon after mounting the crystals. We intend to do further cryogenic gluing and coating tests at a later date when the instrument is opened again and the crystals are going to be unmounted.
VI Conclusion
We have shown that the KDP is an excellent thermal blocking filter and practical to use in an astronomical context, with limitation on its operating temperature, and birefringence. KDP offers good optical characteristics, and can strongly reduce atmospheric or instrumental thermal background. We have demonstrated the use of a warm narrow band filter in conjunction with a 2.5 cutoff detector, which has been enabled by the high thermal IR absorbance of the KDP. While we cannot demonstrate a photometric gain in the warm narrowband over the J band, we measure the night sky surface brightness within the bandpass, and find it to be 18.5 mag arcsec-2. KDP has the potential to become a mainstay in the SWIR astronomy as a heat absorbing, thermal emission blocking, shortpass filter material. It offers a way to improve the SWIR broad- and especially narrowband filter efficiency, and off-band blocking characteristics, can be used for enhancing performance of existing instrumentation suffering from thermal leaks, and has potential for realizing new instrument concepts.
@@ VI Conclusion
MIA conceptualized the project. JPUF acquired hardware funding for the project. JKMV managed the project. JKMV, AAD, and SA recorded the observatory site data. JKMV and AAD processed and analyzed the data. ANS and MIA carried out the thermal analysis. DK and PS recorded the optical density data, and designed and manufactured the NB filter. JKMV wrote the original draft. All authors contributed equally to the review and editing.